# Quantifying $S_{8}$ tension and evidence for interacting dark energy from redshift-space distortion measurements

### v1

@@ I Introduction
Nowadays it is widely accepted that an exotic source of energy with negative pressure known as Dark Energy (DE) is responsible for the accelerated expansion of the universe Riess et al. (1998); Perlmutter et al. (1999) and somehow could be linked with its structure formation history. [^1] Alternatively, modifications to general relativity can also be considered to explain the late-time acceleration of the universe. See Ishak (2018); Saridakis et al. (2023); Heisenberg (2019) for a review. In the standard model of cosmology ($\Lambda$CDM), DE is modeled as a cosmological constant ($\Lambda$), independent of the Cold Dark Matter (CDM) content. While the $\Lambda$CDM model can almost perfectly reproduce a large number of observations Aghanim et al. (2020a, b, c); Mossa et al. (2020); Aiola et al. (2020); Alam et al. (2021); Balkenhol et al. (2023), the increasing precision of observational data has revealed some discrepancies or “tensions" at the core of the $\Lambda$CDM framework. Specifically, there is a tension at more than $5\sigma$ between the early and late-time estimates of the present-day expansion rate of the universe, known as the Hubble constant ($H_{0}$) tension Riess et al. (2022); Verde et al. (2019); Knox and Millea (2020); Di Valentino et al. (2021a, b); Di Valentino (2022); Kamionkowski and Riess (2023); Verde et al. (2023); Di Valentino and Brout (2024); Breuval et al. (2024); Li et al. (2024a); Murakami et al. (2023). Likewise, cosmic shear surveys and Planck-2018 Cosmic Microwave Background (CMB) anisotropy measurements point towards yet another tension surrounding the value of the weighted amplitude of matter fluctuations $S_{8}=\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$ (where $\sigma_{8}$ is the amplitude of matter fluctuations on scales of $8h^{-1},\mathrm{Mpc}$ and $\Omega_{\rm m}$ is the present-day density parameter of matter) as inferred within $\Lambda$CDM, see e.g. Refs. Abbott et al. (2022); Di Valentino et al. (2021c); Di Valentino and Bridle (2018); Abbott et al. (2023); Tröster et al. (2020); Heymans et al. (2021); Dalal et al. (2023); Chen et al. (2024a); Kim et al. (2024); Faga et al. (2024); Harnois-Deraps et al. (2024); Dvornik et al. (2023). This is known as the $S_{8}$ tension. These two tensions have drawn the community’s attention to investigate the extensions of the $\Lambda$CDM model which can explain these discrepancies (see the reviews Di Valentino et al. (2021b); Abdalla et al. (2022); Perivolaropoulos and Skara (2022); Khalife et al. (2024); Di Valentino and Brout (2024) and references therein).
@@
The tension in $S_{8}$ is directly related to the formation and evolution of cosmic structures. The CMB anisotropy measurements from Planck-2018 in the $\Lambda$CDM model provide the best fit: $S_{8}=0.834\pm 0.016$ Aghanim et al. (2020b). This estimate finds an increasing statistical tension of $2$-$3\sigma$ levels with cosmic shear measurements Di Valentino et al. (2021c). For Weak Gravitational Lensing (WGL), the Kilo-Degree Survey (KiDS-1000) analysis provided the constraint: $S_{8}=0.759^{+0.024}_{-0.021}$ Asgari et al. (2021), which is in a $3\sigma$ tension with Planck-2018; the Hyper Suprime-Cam Year 3 reported a $2\sigma$ tension ($S_{8}=0.776^{+0.032}_{-0.033}$alal et al. (2023), and the most recent Dark Energy Survey Year 3 (DES-Y3) analyses measured $S_{8}=0.759^{+0.025}_{-0.023}$ Amon et al. (2022) with a $2.3\sigma$. It is worth mentioning that the analysis of galaxy clustering and WGL of the DES-Y3 data combining three two-point functions (3 $\times$ 2pt analysis) and galaxy positions (with an improvement in signal-to-noise relative to DES-Y1 by a factor of 2.1), gives $S_{8}=0.776^{+0.017}_{-0.017}$ Abbott et al. (2022) in a flat $\Lambda$CDM model. On the other hand, the Redshift Space Distortions (RSD) data exhibit a $2.2\sigma$ tension ($S_{8}=0.762^{+0.030}_{-0.025}$) with Planck-2018 Nunes and Vagnozzi (2021). Combining WGL, real-space clustering, and RSD data in a modified gravity context, the authors in Skara and Perivolaropoulos (2020) found an increase in the growth tension: from $3.5\sigma$ (considering only $f\sigma_{8}$ data) to $6\sigma$ (taking into account also the $E_{g}$ data).
@@
A somewhat crucial difference between the $H_{0}$-tension and the $S_{8}$-tension is that, while for $H_{0}$ we can compare direct measurements in the late universe (which are notably independent of the cosmological model) with model-dependent estimates inferred from early universe observations, all measurements of $S_{8}$ are intrinsically model-dependent. Therefore, even assuming the same $\Lambda$CDM model, early universe measurements with the CMB and late universe measurements with LSS are in disagreement. Numerous approaches have been proposed to address the $S_{8}$ tension, such as decaying dark matter Di Valentino et al. (2018); Franco Abellán et al. (2022); He et al. (2023); Tanimura et al. (2023); Fuß et al. (2024), active and sterile neutrinos Battye and Moss (2014); Feng et al. (2017); Caramete and Popa (2014); Feng et al. (2020); Camarena and Cyr-Racine (2024), modified gravity models Nesseris et al. (2017); Marra and Perivolaropoulos (2021); Acero and Oliveros (2024); de Araujo et al. (2021); Archidiacono et al. (2022); Uzan and Pitrou (2024); Adi and Kovetz (2021); De Felice and Mukohyama (2021); da Silva and Velten (2024); Luongo et al. (2023); Nguyen et al. (2023), among many others Meerburg (2014); Hlozek et al. (2015); Camera et al. (2019); Di Valentino et al. (2020a); Moreno-Pulido and Sola (2020); Lucca (2021); Basilakos et al. (2024); Toda et al. (2024); Akarsu et al. (2024a); Terasawa et al. (2024); Akarsu et al. (2024b); Carrion et al. (2024); Cheek et al. (2024); Chen et al. (2024b); Chakraborty et al. (2024); Stahl et al. (2024); Amon and Efstathiou (2022). A model that has been extensively studied in the literature and shown remarkable promise in resolving various issues currently associated with the standard model is the so-called Interacting Dark Energy (IDE) Kumar and Nunes (2016); Murgia et al. (2016); Kumar and Nunes (2017); Di Valentino et al. (2017); Yang et al. (2020); Kumar (2021); Forconi et al. (2024); Benisty et al. (2024); Pourtsidou and Tram (2016); Di Valentino (2021); Di Valentino and Mena (2020); Nunes and Di Valentino (2021); Yang et al. (2018); von Marttens et al. (2020); Lucca and Hooper (2020); Zhai et al. (2023a); Bernui et al. (2023); Hoerning et al. (2023); Giarè et al. (2024a); Escamilla et al. (2023); van der Westhuizen and Abebe (2024); Silva et al. (2024); Di Valentino et al. (2020b); Li et al. (2024b); Pooya (2024); Halder et al. (2024); Castello et al. (2024); Pan and Yang (2023); Yao and Meng (2023); Mishra et al. (2023); Nunes et al. (2016), where a non-gravitational interaction between DE and DM is postulated (see e.g. Ref. Wang et al. (2024) for a recent review). Recently, in Ref. Giarè et al. (2024b) it was shown by some of us that the new Baryon Acoustic Oscillations (BAO) measurements from the Dark Energy Spectroscopic Instrument (DESI), in conjunction with Planck-2018 CMB data, exhibit a preference for DE-DM interaction over $\Lambda$CDM while reducing the $H_{0}$-tension. In this paper, we explore in detail the role of RSD samples in this cosmological context, considering both scenarios where DM can transfer energy to DE and the opposite case. Our analysis reveals new perspectives, establishing new upper bounds for the case $\xi<0$ and providing evidence in favor of $\xi>0$ when RSD data are considered. Additionally, we demonstrate that models with $\xi>0$ could potentially alleviate the tension observed in $S_{8}$.
@@ III Datasets and methodology
  • Redshift Space Distortions (RSD): 22 measurements of $f\sigma_{8}(z)$ spanning the redshift range $1.9440.02<z<$ 1.944 as gathered from various surveys and sources in the literature and summarized in Table I of Ref Sagredo et al. (2018). Note that different $f\sigma_{8}(z)$ measurements can involve different assumptions and be subject to distinct uncertainties. Therefore, it is essential to assess their internal consistency. The internal consistency of the RSD measurements adopted in this study has been assessed using a Bayesian model comparison framework, as detailed in Ref.Sagredo et al. (2018).
@@
Parameter RSD DESI+PP RSD+DESI+PP RSD+DESI+PP+CC
$\xi$ $>-0.0192$ ($>-0.0331$) $>-0.0729$ ($>-0.177$) $>-0.00568$ ($>-0.0133$) $>-0.00563$ ($>-0.0133$)
$\Omega{}_{\rm m}$ $0.218^{+0.036}_{-0.067}$ ($0.218^{+0.11}_{-0.090}$) $0.297^{+0.016}_{-0.013}$ ($0.297^{+0.030}_{-0.031}$) $0.2822\pm 0.0075$ ($0.282^{+0.015}_{-0.015}$) $0.2826\pm 0.0075$ ($0.283^{+0.014}_{-0.015}$)
$S_{8}$ $0.670^{+0.058}_{-0.051}$ ($0.67^{+0.10}_{-0.11}$) - $0.773\pm 0.025$ ($0.773^{+0.049}_{-0.048}$) $0.774\pm 0.024$ ($0.774^{+0.048}_{-0.047}$)
$H_{0}$ (km/s/Mpc) $78^{+10}_{-10}$ ($78^{+20}_{-20}$) $70.17^{+0.82}_{-1.0}$ ($70.2^{+1.9}_{-1.8}$) $68.79\pm 0.66$ ($68.8^{+1.3}_{-1.3}$) $68.79\pm 0.65$ ($68.8^{+1.3}_{-1.3}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $-0.0004$ $-0.294$ $-0.12$ $1.438$
@@
Parameter RSD DESI+PP RSD+DESI+PP RSD+DESI+PP+CC
$\xi$ $<0.630$ ($<0.949$) $0.26^{+0.12}_{-0.15}$ ($0.26^{+0.23}_{-0.25}$) $0.044^{+0.017}_{-0.020}$ ($0.044^{+0.036}_{-0.036}$) $0.044^{+0.016}_{-0.020}$ ($0.044^{+0.036}_{-0.035}$)
$\Omega_{\rm m}$ $0.573^{+0.19}_{-0.086}$ ($0.57^{+0.20}_{-0.28}$) $0.356^{+0.024}_{-0.029}$ ($0.356^{+0.051}_{-0.050}$) $0.312\pm 0.013$ ($0.312^{+0.027}_{-0.025}$) $0.313^{+0.012}_{-0.014}$ ($0.313^{+0.027}_{-0.024}$)
$S_{8}$ $1.03^{+0.20}_{-0.13}$ ($1.03^{+0.25}_{-0.29}$) - $0.823\pm 0.031$ ($0.823^{+0.061}_{-0.059}$) $0.824\pm 0.030$ ($0.824^{+0.061}_{-0.058}$)
$H_{0}$ (km/s/Mpc) $56.7^{+3.2}_{-6.3}$ ($57^{+10}_{-10}$) $67.0^{+1.6}_{-1.4}$ ($67.0^{+2.8}_{-3.0}$) $68.99\pm 0.68$ ($69.0^{+1.3}_{-1.3}$) $68.97\pm 0.67$ ($69.0^{+1.3}_{-1.3}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $0.25$ $-3.448$ $0.18$ $-2.80$
@@ IV Results and Discussions
$\Omega_{\rm b}h^{2}$ $0.02238\pm 0.00014$ ($0.02238^{+0.00028}_{-0.00028}$) $0.02238\pm 0.00014$ ($0.02238^{+0.00028}_{-0.00029}$) $0.02254\pm 0.00013$ ($0.02254^{+0.00026}_{-0.00026}$) $0.02250\pm 0.00013$ ($0.02250^{+0.00026}_{-0.00026}$)
$\Omega_{\rm c}h^{2}$ $0.1350^{+0.0072}_{-0.013}$ ($0.135^{+0.020}_{-0.016}$) $0.1236\pm 0.0022$ ($0.1236^{+0.0046}_{-0.0042}$) $0.1207^{+0.0016}_{-0.0018}$ ($0.1207^{+0.0033}_{-0.0032}$) $0.1214\pm 0.0017$ ($0.1214^{+0.0033}_{-0.0031}$)
$100\theta_{\mathrm{s}}$ $1.04187\pm 0.00029$ ($1.04187^{+0.00056}_{-0.00056}$) $1.04188\pm 0.00030$ ($1.04188^{+0.00057}_{-0.00061}$) $1.04208\pm 0.00028$ ($1.04208^{+0.00055}_{-0.00057}$) $1.04205\pm 0.00028$ ($1.04205^{+0.00057}_{-0.00053}$)
$\tau_{\mathrm{reio}}$ $0.0544\pm 0.0076$ ($0.054^{+0.016}_{-0.014}$) $0.0545\pm 0.0072$ ($0.054^{+0.014}_{-0.015}$) $0.0598^{+0.0071}_{-0.0080}$ ($0.060^{+0.016}_{-0.014}$) $0.0588^{+0.0068}_{-0.0076}$ ($0.059^{+0.015}_{-0.014}$)
$n_{\mathrm{s}}$ $0.9651\pm 0.0040$ ($0.9651^{+0.0080}_{-0.0080}$) $0.9657\pm 0.0042$ ($0.9657^{+0.0085}_{-0.0082}$) $0.9705\pm 0.0036$ ($0.9705^{+0.0071}_{-0.0072}$) $0.9697\pm 0.0036$ ($0.9697^{+0.0070}_{-0.0070}$)
$\log(10^{10}A_{\mathrm{s}})$ $3.046\pm 0.015$ ($3.046^{+0.032}_{-0.028}$) $3.046\pm 0.014$ ($3.046^{+0.028}_{-0.028}$) $3.052^{+0.014}_{-0.016}$ ($3.052^{+0.030}_{-0.028}$) $3.051^{+0.013}_{-0.015}$ ($3.051^{+0.030}_{-0.028}$)
$\xi$ $<0.205$ ($<0.388$) $0.035^{+0.014}_{-0.017}$ ($0.035^{+0.030}_{-0.031}$) $0.024^{+0.011}_{-0.015}$ ($0.024^{+0.022}_{-0.024}$) $0.029^{+0.012}_{-0.015}$ ($0.029^{+0.025}_{-0.026}$)
$H_{0}$ (km/s/Mpc) $65.8^{+1.5}_{-0.99}$ ($65.8^{+2.1}_{-2.6}$) $67.07\pm 0.60$ ($67.1^{+1.2}_{-1.2}$) $68.09\pm 0.44$ ($68.09^{+0.89}_{-0.86}$) $67.88\pm 0.41$ ($67.88^{+0.80}_{-0.82}$)
$\Omega_{\rm m}$ $0.366^{+0.025}_{-0.048}$ ($0.366^{+0.076}_{-0.059}$) $0.3262^{+0.0098}_{-0.011}$ ($0.326^{+0.021}_{-0.019}$) $0.3104\pm 0.0072$ ($0.310^{+0.015}_{-0.013}$) $0.3138\pm 0.0069$ ($0.314^{+0.014}_{-0.013}$)
$S_{8}$ $0.787\pm 0.024$ ($0.787^{+0.044}_{-0.046}$) $0.809\pm 0.012$ ($0.809^{+0.023}_{-0.022}$) $0.7926^{+0.0087}_{-0.010}$ ($0.793^{+0.020}_{-0.018}$) $0.7950\pm 0.0091$ ($0.795^{+0.018}_{-0.018}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $0.9$ $-2.96$ $-1.22$ $-4.5$

### v2

# Quantifying the $S_{8}$ tension and evidence for interacting dark energy from redshift-space distortion measurements

@@ I Introduction
Nowadays it is widely accepted that an exotic source of energy with negative pressure known as Dark Energy (DE) is responsible for the accelerated expansion of the universe Riess et al. (1998); Perlmutter et al. (1999) and somehow could be linked with its structure formation history. [^1] Alternatively, modifications to general relativity can also be considered to explain the late-time acceleration of the universe. See Ishak (2018); Saridakis et al. (2023); Heisenberg (2019) for a review. In the standard model of cosmology ($\Lambda$CDM), DE is modeled as a cosmological constant ($\Lambda$), independent of the Cold Dark Matter (CDM) content. While the $\Lambda$CDM model can almost perfectly reproduce a large number of observations Aghanim et al. (2020a, b, c); Mossa et al. (2020); Aiola et al. (2020); Alam et al. (2021); Balkenhol et al. (2023), the increasing precision of observational data has revealed some discrepancies or “tensions" at the core of the $\Lambda$CDM framework. Specifically, there is a tension at more than $5\sigma$ between the early and late-time estimates of the present-day expansion rate of the universe, known as the Hubble constant ($H_{0}$) tension Riess et al. (2022); Verde et al. (2019); Knox and Millea (2020); Di Valentino et al. (2021a, b); Di Valentino (2022); Kamionkowski and Riess (2023); Verde et al. (2023); Di Valentino and Brout (2024); Breuval et al. (2024); Li et al. (2024a); Murakami et al. (2023). Likewise, cosmic shear surveys and Planck-2018 Cosmic Microwave Background (CMB) anisotropy measurements point towards yet another tension surrounding the value of the weighted amplitude of matter fluctuations $S_{8}=\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$ (where $\sigma_{8}$ is the amplitude of matter fluctuations on scales of $8h^{-1}\mathrm{Mpc}$, and $\Omega_{\rm m}$ is the present-day density parameter of matter) as inferred within $\Lambda$CDM, see e.g. Refs. Abbott et al. (2022); Di Valentino et al. (2021c); Di Valentino and Bridle (2018); Abbott et al. (2023); Tröster et al. (2020); Heymans et al. (2021); Dalal et al. (2023); Chen et al. (2024a); Kim et al. (2024); Faga et al. (2024); Harnois-Deraps et al. (2024); Dvornik et al. (2023). This is known as the $S_{8}$ tension. These two tensions have drawn the community’s attention to investigate the extensions of the $\Lambda$CDM model which can explain these discrepancies (see the reviews Di Valentino et al. (2021b); Abdalla et al. (2022); Perivolaropoulos and Skara (2022a); Khalife et al. (2024); Di Valentino and Brout (2024) and references therein).
@@
The tension in $S_{8}$ is directly related to the formation and evolution of cosmic structures. The CMB anisotropy measurements from Planck-2018 in the $\Lambda$CDM model provide the best fit: $S_{8}=0.834\pm 0.016$ Aghanim et al. (2020b). This estimate finds an increasing statistical tension of $2$-$3\sigma$ levels with cosmic shear measurements Di Valentino et al. (2021c). For Weak Gravitational Lensing (WGL), the Kilo-Degree Survey (KiDS-1000) analysis provided the constraint: $S_{8}=0.759^{+0.024}_{-0.021}$ Asgari et al. (2021), which is in a $3\sigma$ tension with Planck-2018; the Hyper Suprime-Cam Year 3 reported a $2\sigma$ tension ($S_{8}=0.776^{+0.032}_{-0.033}$alal et al. (2023), and the most recent Dark Energy Survey Year 3 (DES-Y3) analyses measured $S_{8}=0.759^{+0.025}_{-0.023}$ Amon et al. (2022) with a $2.3\sigma$. It is worth mentioning that the analysis of galaxy clustering and WGL of the DES-Y3 data combining three two-point functions (3 $\times$ 2pt analysis) and galaxy positions (with an improvement in signal-to-noise relative to DES-Y1 by a factor of 2.1), gives $S_{8}=0.776^{+0.017}_{-0.017}$ Abbott et al. (2022) in a flat $\Lambda$CDM model. On the other hand, Redshift Space Distortion (RSD) data alone show a $3.1\sigma$ tension with the Planck-CMB results Nunes and Vagnozzi (2021); Kazantzidis and Perivolaropoulos (2018). When combining weak gravitational lensing, real-space clustering, and RSD data, the authors of Skara and Perivolaropoulos (2020) observed a significant increase in the tension related to the growth of structure, rising from $3.5\sigma$ (based on RSD data alone) to $6\sigma$ when including $E_{g}$ data as well. See Perivolaropoulos and Skara (2022b) for a more discussion on this topic.
@@
A somewhat crucial difference between the $H_{0}$-tension and the $S_{8}$-tension is that, while for $H_{0}$ we can compare direct measurements in the late universe (which are notably independent of the cosmological model) with model-dependent estimates inferred from early universe observations, all measurements of $S_{8}$ are intrinsically model-dependent. Therefore, even assuming the same $\Lambda$CDM model, early universe measurements with the CMB and late universe measurements with LSS are in disagreement. Numerous approaches have been proposed to address the $S_{8}$ tension, such as decaying dark matter Di Valentino et al. (2018); Franco Abellán et al. (2022); He et al. (2023); Tanimura et al. (2023); Fuß et al. (2024), active and sterile neutrinos Battye and Moss (2014); Feng et al. (2017); Caramete and Popa (2014); Feng et al. (2020); Camarena and Cyr-Racine (2024), modified gravity models Nesseris et al. (2017); Marra and Perivolaropoulos (2021); Acero and Oliveros (2024); de Araujo et al. (2021); Archidiacono et al. (2022); Uzan and Pitrou (2024); Adi and Kovetz (2021); De Felice and Mukohyama (2021); da Silva and Velten (2024); Luongo et al. (2023); Nguyen et al. (2023), among many others Meerburg (2014); Hlozek et al. (2015); Camera et al. (2019); Di Valentino et al. (2020a); Moreno-Pulido and Sola (2020); Lucca (2021); Cheek et al. (2023); Basilakos et al. (2024); Toda et al. (2024); Akarsu et al. (2024a); Terasawa et al. (2024); Akarsu et al. (2024b); Carrion et al. (2024); Cheek et al. (2024); Chen et al. (2024b); Chakraborty et al. (2024); Stahl et al. (2024); Amon and Efstathiou (2022). A model that has been extensively studied in the literature and shown remarkable promise in resolving various issues currently associated with the standard model is the so-called Interacting Dark Energy (IDE) Kumar and Nunes (2016); Murgia et al. (2016); Kumar and Nunes (2017); Di Valentino et al. (2017); Yang et al. (2020); Kumar (2021); Forconi et al. (2024); Benisty et al. (2024); Pourtsidou and Tram (2016); Di Valentino (2021); Di Valentino and Mena (2020); Nunes and Di Valentino (2021); Yang et al. (2018); von Marttens et al. (2020); Lucca and Hooper (2020); Zhai et al. (2023a); Bernui et al. (2023); Hoerning et al. (2023); Giarè et al. (2024a); Escamilla et al. (2023); van der Westhuizen and Abebe (2024); Silva et al. (2024); Di Valentino et al. (2020b); Li et al. (2024b); Pooya (2024); Halder et al. (2024); Castello et al. (2024); Pan and Yang (2023); Yao and Meng (2023); Mishra et al. (2023); Nunes et al. (2016), where a non-gravitational interaction between DE and DM is postulated (see e.g. Ref. Wang et al. (2024) for a recent review). Recently, in Ref. Giarè et al. (2024b) it was shown by some of us that the new Baryon Acoustic Oscillations (BAO) measurements from the Dark Energy Spectroscopic Instrument (DESI), in conjunction with Planck-2018 CMB data, exhibit a preference for DE-DM interaction over $\Lambda$CDM while reducing the $H_{0}$-tension. In this paper, we explore in detail the role of RSD samples in this cosmological context, considering both scenarios where DM can transfer energy to DE and the opposite case. Our analysis offers fresh insights, setting new upper bounds on the interaction parameter in both cases—$\xi<0$ (DM to DE) and $\xi>0$ (DE to DM)—when RSD data are incorporated Additionally, we demonstrate that models with $\xi>0$ could potentially alleviate the tension observed in $S_{8}$.
@@ III Datasets and methodology
  • Redshift Space Distortions (RSD): We compile 20 measurements of $f\sigma_{8}(z)$ spanning the redshift range $1.9440.02<z<$ 1.944, sourced from various surveys and references in the literature, as summarized in Avila et al. (2022). The data points are selected as per the following criteria:
    • We include $f\sigma_{8}(z)$ measurements from uncorrelated redshift bins when they are based on the same cosmological tracer. Data from potentially correlated redshift bins are also considered, but they correspond to different cosmological tracers.
    • Only direct measurements of $f\sigma_{8}(z)$ are included.
    • For surveys with multiple $f\sigma_{8}(z)$ measurements across different data releases, only the most recent measurements are considered.
    The compilation above may include fewer data points than typically assumed in recent literature, but it ensures that our results remain statistically consistent by adhering to a rigorous selection criteria for RSD data.
@@
Parameter RSD DESI+PP DESI+PP+RSD DESI+PP+RSD+CC
$\xi$ $\xi>-0.0250$ ($\xi>-0.0440$) $\xi>-0.0729$ ($\xi>-0.177$) $\xi>-0.0200$ ($\xi>-0.0430$) $\xi>-0.0197$ ($\xi>-0.0422$)
$\Omega_{\mathrm{m}}$ $0.253^{+0.053}_{-0.074}$ ($0.25^{+0.13}_{-0.12}$) $0.297^{+0.016}_{-0.013}$ ($0.297^{+0.030}_{-0.031}$) $0.304\pm 0.012$ ($0.304^{+0.024}_{-0.023}$) $0.304\pm 0.012$ ($0.304^{+0.023}_{-0.022}$)
$S_{8}$ $0.738\pm 0.068$ ($0.74\pm 0.14$) -- $0.775\pm 0.027$ ($0.775^{+0.052}_{-0.053}$) $0.775\pm 0.026$ ($0.775^{+0.052}_{-0.052}$)
$H_{0}\mathrm{(km/s/Mpc)}$- $70.17^{+0.82}_{-1.0}$ ($70.2^{+1.9}_{-1.8}$) $69.73\pm 0.78$ ($69.7^{+1.5}_{-1.5}$) $69.67\pm 0.75$ ($69.7^{+1.5}_{-1.4}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $0.031$ $-0.0009$ $-0.042$ $0.048$
@@
Parameter RSD DESI+PP DESI+PP+RSD DESI+PP+RSD+CC
$\xi$ $\xi<0.575$ ($\xi<0.926$) $0.26^{+0.12}_{-0.15}$ ($0.26^{+0.23}_{-0.25}$) $\xi<0.0396$ ($\xi<0.0717$) $\xi<0.0396$ ($\xi<0.0716$)
$\Omega_{\mathrm{m}}$ $0.582^{+0.18}_{-0.083}$ ($0.58^{+0.20}_{-0.28}$) $0.356^{+0.024}_{-0.029}$ ($0.356^{+0.051}_{-0.050}$) $0.317\pm 0.013$ ($0.317^{+0.026}_{-0.024}$) $0.316^{+0.012}_{-0.013}$ ($0.316^{+0.026}_{-0.024}$)
$S_{8}$ $1.08^{+0.19}_{-0.11}$ ($1.08^{+0.24}_{-0.28}$) -- $0.812^{+0.027}_{-0.031}$ ($0.812^{+0.060}_{-0.054}$) $0.812^{+0.028}_{-0.032}$ ($0.812^{+0.061}_{-0.057}$)
$H_{0}\mathrm{(km/s/Mpc)}$- $67.0^{+1.6}_{-1.4}$ ($67.0^{+2.8}_{-3.0}$) $69.37\pm 0.78$ ($69.4^{+1.6}_{-1.5}$) $69.31\pm 0.77$ ($69.3^{+1.5}_{-1.5}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $-0.15$ $-3.45$ $-0.37$ $-0.26$
@@ IV Results and Discussions
$\Omega_{\rm b}h^{2}$ $0.02238\pm 0.00014$ ($0.02238^{+0.00028}_{-0.00028}$) $0.002237\pm 0.00015$ ($0.002237^{+0.00029}_{-0.00028}$) $0.002253\pm 0.00013$ ($0.002253^{+0.00029}_{-0.00027}$) $0.002250\pm 0.00013$ ($0.002250^{+0.00027}_{-0.00026}$)
$\Omega_{\rm c}h^{2}$ $0.1350^{+0.0072}_{-0.013}$ ($0.135^{+0.020}_{-0.016}$) $0.1227^{+0.0018}_{-0.0024}$ ($0.1227^{+0.0044}_{-0.0040}$) $0.1199^{+0.0013}_{-0.0019}$ ($0.1199^{+0.0033}_{-0.0029}$) $0.1207^{+0.0014}_{-0.0019}$ ($0.1207^{+0.0033}_{-0.0030}$)
$100\theta_{\mathrm{s}}$ $1.04187\pm 0.00029$ ($1.04187^{+0.00056}_{-0.00056}$) $1.04189\pm 0.00030$ ($1.04189^{+0.00060}_{-0.00055}$) $1.04206\pm 0.00028$ ($1.04206^{+0.00055}_{-0.00057}$) $1.04204\pm 0.00028$ ($1.04204^{+0.00056}_{-0.00053}$)
$\tau_{\mathrm{reio}}$ $0.0544\pm 0.0076$ ($0.054^{+0.016}_{-0.014}$) $0.0543^{+0.0065}_{-0.0077}$ ($0.054^{+0.015}_{-0.013}$) $0.0603^{+0.0069}_{-0.0082}$ ($0.060^{+0.016}_{-0.014}$) $0.0590\pm 0.0071$ ($0.059^{+0.015}_{-0.013}$)
$n_{\mathrm{s}}$ $0.9651\pm 0.0040$ ($0.9651^{+0.0080}_{-0.0080}$) $0.9655\pm 0.0041$ ($0.9655^{+0.0082}_{-0.0078}$) $0.9705\pm 0.0037$ ($0.9705^{+0.0075}_{-0.0073}$) $0.9695\pm 0.0035$ ($0.9695^{+0.0071}_{-0.0067}$)
$\log(10^{10}A_{\mathrm{s}})$ $3.046\pm 0.015$ ($3.046^{+0.032}_{-0.028}$) $3.045\pm 0.014$ ($3.045^{+0.029}_{-0.027}$) $3.053^{+0.014}_{-0.016}$ ($3.053^{+0.031}_{-0.029}$) $3.051\pm 0.014$ ($3.051^{+0.028}_{-0.026}$)
$\xi$ $<0.205$ ($<0.388$) $0.0250^{+0.0079}_{-0.023}$ ($<0.0545$) $<0.0214$ ($<0.0405$) $<0.0259$ ($<0.0449$)
$H_{0}$ (km/s/Mpc) $65.8^{+1.5}_{-0.99}$ ($65.8^{+2.1}_{-2.6}$) $67.11\pm 0.58$ ($67.1^{+1.2}_{-1.1}$) $68.13\pm 0.43$ ($68.13^{+0.83}_{-0.82}$) $67.92\pm 0.41$ ($67.92^{+0.81}_{-0.77}$)
$\Omega_{\rm m}$ $0.366^{+0.025}_{-0.048}$ ($0.366^{+0.076}_{-0.059}$) $0.3238^{+0.0088}_{-0.010}$ ($0.324^{+0.020}_{-0.018}$) $0.3084^{+0.0060}_{-0.0072}$ ($0.308^{+0.014}_{-0.012}$) $0.3118^{+0.0062}_{-0.0071}$ ($0.312\pm 0.013$)
$S_{8}$ $0.787\pm 0.024$ ($0.787^{+0.044}_{-0.046}$) $0.813\pm 0.012$ ($0.813^{+0.024}_{-0.024}$) $0.796\pm 0.010$ ($0.796\pm 0.019$) $0.798\pm 0.010$ ($0.798^{+0.019}_{-0.019}$)
$\Delta\chi^{2}_{\mathrm{min}}$ $0.9$ $-1.98$ $-1.13$ $-0.76$